OpenStudy (anonymous):

Find the equations of the lines tangent to y=x^2 and pass through the point (1,-3).

OpenStudy (anonymous):

What I've already gotten it down to is [(x^2-3)/(x-1)]=2x

OpenStudy (anonymous):

Sorry, x^2+3

OpenStudy (anonymous):

I figured it out. \[(x^2+3)/(x-1)=2x\] \[2x^2-2x-x^2-3=0\] \[(x-3)(x+1)=0\] \[x=3, -1\]

OpenStudy (anonymous):

\[y+3=-2(x-1)\] \[y=-2x-1\]

OpenStudy (anonymous):

\[y+3=6(x-1)\] \[y=6x-9\]